OpenStudy (anonymous):

Solve using the guess-and-test strategy. The math club has 57 members who are in sixth or seventh grades. There are twice as many sixth graders as seventh graders. How many sixth graders are in the club? A. 38 B. 37 C. 19 D. 18 plz help meh if u know the answer to this question

OpenStudy (raden):

let the number of members of sixth graders = x, so number of members of seventh graders = 2x according information above, the expression algebraic for this problem : x+2x = 57 3x = 57 x = 57/3 = 19 therefore, there are 19 members of sixth graders